# Constant-Overhead Magic State Distillation

### v1

@@ 1.1 Overview of Our Results
Since we have $k,d=\Theta(n)$, $\gamma$ in such a concatenated protocol would be made arbitrarily close to 00. However, we note that this could be achieved with any constant rate $k/n$ and any $d$ which diverges to infinity as $n\to\infty$. In our case, it turns out that because the parameters are asymptotically good, a stronger claim is possible. Using just one round of the magic state distillation subroutine, i.e., without concatenating it with itself, distillation with $\gamma=0$ is possible. Essentially, this is true because the error suppression is strong enough (following from $d=\Theta(n)$) that one round of magic state distillation, with a large enough $n$, is enough to make the target error rate as small as desired. Because $k=\Theta(n)$, this means that the number of outputted |$\ket{U_{10}}$ states grows in proportion to the number of input states; i.e., the overhead is constant $\mathcal{O}(1)$ as $\epsilon\to 0$. When combined with the conversion to the magic states |$\ket{CCZ}$ (with only a constant-factor loss), this establishes Theorem 1.1. This argument on the error analysis will be made precise in Section 6.
@@ 1.2 Discussion and Future Directions
As discussed, the protocols we construct naturally distil a $10$-qubit magic state for a $10$-qubit non-Clifford gate, but this magic state may be converted to and from |$\ket{CCZ}$ by stabiliser operations; thus, our protocols may distil the |$\ket{CCZ}$ state. Distillation of |$\ket{T}$ states requires further conversions [20, 21, 22, 23]. Algebraic geometry codes seem to provide the desired framework to construct matrices satisfying polynomial equations like Definition 1.1.1. The question one might ask, therefore, is whether magic states for other gates can be distilled using a similar framework, i.e., similar polynomial relations to achieve transversality. Ideally, these might allow for constructions of codes for fields smaller than $\mathbb{F}_{1024}$ used for our construction.
Open Question 1: Can we construct asymptotically good quantum codes supporting other transversally implementable non-Clifford gates than ours? Can $s<10$ be achieved?
@@ 2.3.3 The Construction of Algebraic Geometry Codes, their Parameters and Decoding
To construct algebraic geometry codes, it is therefore necessary to have function fields $F$ of genus $g$ with a large number of rational places compared to their genus. Indeed, to establish a non-zero distance on the dual code $C_{\mathcal{L}}(D,A)^{\perp}$ (which will be crucial for our purpose, magic state distillation), we need $2g-1<n$, where $n$ is a number of rational places used to build the code. Given that the Hasse-Weil bound [39, Theorem 5.2.3] says that the number $N(F)$ of rational places of a function field $F/\mathbb{F}_{q}$ of genus $g$ always satisfies
@@
then there is an efficient classical algorithm to decode the (classical) code $C_{\Omega}(D,A)$ from errors of weight at most $t$, as presented by [39, Decoding Algorithm 8.5.4]. Note that if the weight of errors exceeds $t$, the decoder may not be able to find the correct codeword, but we assume that the decoder, even in such a case, still outputs some codeword that may have a logical error; that is, we work on a setting without post-selection. The runtime of this decoder is $O(\mathrm{poly}(n))$, where the dominant step is to solve a system of linear equations as presented in Section 8.5 of [39].
@@ 4.1 Summary of Magic State Distillation Protocols
    The outputted $k$ |$\ket{M}$ states, each of which are states of one qudit, may equally well be thought of as $k$ copies of the $s$-qubit state |$\ket{U_{s}}$.
@@
Readers familiar with [13] and [19] will note the similarity of step 3 to those procedures. Those works construct $[[n,k,d]]$ quantum codes (on qubits in the first case, and on prime-dimensional qudits in the second case) which support transversal gates that are diagonal, non-Clifford and in the third level of the Clifford hierarchy. Calling this non-Clifford gate $U$ for convenience, they each make use of a distillation subroutine, which goes as follows, on either qubits, or prime-dimensional qudits.
@@
    Prepare $k$ qubits/qudits in the state |$\ket{+}^{\otimes k}$, where $\ket{+}$ is an equal superposition of all computational basis states. Encode these qubit/qudits into the quantum code.
@@
Then, [13] and [19] concatenate this distillation subroutine to make a larger distillation routine that may arbitrarily suppress the error rate. In particular, the distillation routine will have multiple levels where the output of one level is used as the input to the next level. Where $[[n,k,d]]$ are the parameters of the quantum code at hand, both papers show that, in order to generate a supply of magic states with error rate below $\epsilon$, the number of input magic states (of some (small) constant error rate) required per output state is
@@
which is called the overhead of the magic state distillation protocol, where
@@ 6 Analysis of Error Suppression and Overhead
In this section, we analyse the error and the overhead of the distillation protocol presented in this paper. We consider the task of magic state distillation under the local stochastic error model described in Section 2.4. As described in Sections 4, our distillation protocol, using $[[n,k,d]]$ triorthogonal codes on $q$-dimensional qudits with $q=p^{s}$, starts with $\zeta=Cn$ noisy $CCZ$ states |$\ket{CCZ}$ at physical error rate $p_{\mathrm{ph}}$, converts them into $n$ noisy magic states |$\ket{M}$ on qudits, distils $k$ magic states |$\ket{M}$ and converts them into $l=k$ $CCZ$ states. In particular, as presented in Section 4, our protocol is designed for $p=2$, $s=10$, $C=O(1)$, $k=\Theta(n)$, and $d=\Theta(n)$. The efficient decoder of the algebraic geometry codes (i.e., of our triorthogonal codes) has the decoding radius of $t=\Theta(n)$ and thus can correct any $Z$-type errors (which are relevant for magic state distillation) up to weight $t$ on the $n$ qudits. Note that our protocol does not involve post-selection since we use the decoder for error correction rather than error detection (see also Section 4).
@@
Under the local stochastic error model, our protocol for magic state distillation has a constant threshold $p_{\mathrm{th}}>0$ such that if the physical error rate of the $\zeta$ initial $CCZ$ states is below the threshold, i.e., $p_{\mathrm{ph}}<p_{\mathrm{th}}$, then the protocol outputs $\xi(\zeta)$ final $CCZ$ states achieving a target error rate $\epsilon$ exponentially suppressed as
@@ Acknowledgements
All authors extend their thanks to N. Rengaswamy for discussions on CSS-T codes. A. Wills thanks The University of Tokyo, in particular the whole group of H. Yamasaki, for their funding of and hospitality during his visit to the institution, which led to this work.

### v2

@@ 1.1 Overview of Our Results
Since we have $k,d=\Theta(n)$, $\gamma$ in such a concatenated protocol would be made arbitrarily close to 00. However, we note that this could be achieved with any constant rate $k/n$ and any $d$ which diverges to infinity as $n\to\infty$. In our case, it turns out that because the parameters are asymptotically good, a stronger claim is possible. Using just one round of the magic state distillation subroutine, i.e., without concatenating it with itself, distillation with $\gamma=0$ is possible. Essentially, this is true because the error suppression is strong enough (following from $d=\Theta(n)$) that one round of magic state distillation, with a large enough $n$, is enough to make the target error rate as small as desired. To achieve $\gamma=0$, it is also crucial that our protocol, using a polynomial-time decoder for algebraic geometry codes, performs error correction without post-selection to keep the overhead small [^4] We note that the idea of using error correction, rather than post-selection, in magic state distillation has appeared before [26, 27]. This idea was originally due to David Poulin [28].. Because $k=\Theta(n)$, this means that the number of outputted |$\ket{U_{10}}$ states grows in proportion to the number of input states; i.e., the overhead is constant $\mathcal{O}(1)$ as $\epsilon\to 0$. When combined with the conversion to the magic states |$\ket{CCZ}$ (with only a constant-factor loss), this establishes Theorem 1.1. This argument on the error analysis will be made precise in Section 6.
Note on related works
Independent works [29, 30] from ours have also presented asymptotically good quantum codes that support transversal non-Clifford gates. Whereas our protocol for magic state distillation uses $2^{10}$-dimensional qudits for our codes, the codes presented in these works support transversal $CCZ$ gates on qubits, addressing one of the open questions in the first version of this paper. These other works use some similar constructions to ours, and some other techniques, which will likely prove to be complementary to those in the present paper.
However, our contribution is to prove the achievability of constant-overhead magic state distillation in addition to the construction of asymptotically good codes. The use of asymptotically good codes in the existing protocols for magic state distillation leads to the polylogarithmic overhead $\mathcal{O}(\log^{\gamma}(1/\epsilon))$ with $\gamma>0$ arbitrarily small, according to the existing analysis summarized in Equations (4) and (11). Without presenting different protocols and analyses, the works [29, 30] only lead to $\gamma\to 0$, i.e., magic state distillation with polylogarithmic overhead of arbitrarily low degree, rather than constant-overhead magic state distillation. By contrast, our error analysis in Section 6 shows that using the asymptotically good codes in the single-round protocol presented in this work, which uses error correction without post-selection, it is possible to achieve the constant overhead $\gamma=0$, where we can also use the asymptotically good codes in Refs. [29, 30] as well as our asymptotically good codes.
@@ 1.2 Discussion and Future Directions
As discussed, the protocols we construct naturally distil a $10$-qubit magic state for a $10$-qubit non-Clifford gate, but this magic state may be converted to and from |$\ket{CCZ}$ by stabiliser operations; thus, our protocols may distil the |$\ket{CCZ}$ state. Distillation of |$\ket{T}$ states requires further conversions [20, 21, 22, 23]. Algebraic geometry codes provide the desired framework to construct asymptotically good codes from matrices satisfying polynomial equations like Definition 1.1.1. However, in this work, we did not optimise constant factors appearing in these codes and protocols for constant-overhead magic state distillation. The question one might ask, therefore, is whether one may realise practical constant-overhead magic state distillation, which may require improving the codes and protocols with more optimised constant factors.
Open Question 1: Can we construct asymptotically good quantum codes supporting transversally implementable non-Clifford gates with more optimised constant factors? In a practical finite-length regime, can the protocols for constant-overhead magic state distillation be advantageous over the conventional polylogarithmic-overhead protocols?
@@ 2.3.3 The Construction of Algebraic Geometry Codes, their Parameters and Decoding
To construct algebraic geometry codes, it is therefore necessary to have function fields $F$ of genus $g$ with a large number of rational places compared to their genus. Indeed, to establish a non-zero distance on the dual code $C_{\mathcal{L}}(D,A)^{\perp}$ (which will be crucial for our purpose, magic state distillation), we need $2g-1<n$, where $n$ is a number of rational places used to build the code. Given that the Hasse-Weil bound [44, Theorem 5.2.3] says that the number $N(F)$ of rational places of a function field $F/\mathbb{F}_{q}$ of genus $g$ always satisfies
@@
then there is an efficient classical algorithm to decode the (classical) code $C_{\Omega}(D,A)$ from errors of weight at most $t$, as presented by [44, Decoding Algorithm 8.5.4]. Note that if the weight of errors exceeds $t$, the decoder may not be able to find the correct codeword, but we assume that the decoder, even in such a case, still outputs some codeword that may have a logical error; that is, we work on a setting without post-selection. The runtime of this decoder is $O(\mathrm{poly}(n))$, where the dominant step is to solve a system of linear equations as presented in Section 8.5 of [44].
@@ 4.1 Summary of Magic State Distillation Protocols
    The outputted $k$ |$\ket{M}$ states, each of which is a state of one qudit, may equally well be thought of as $k$ copies of the $s$-qubit state |$\ket{U_{s}}$.
@@
Readers familiar with [13] and [19] will note the similarity of step 3 to those procedures. Those works construct $[[n,k,d]]$ quantum codes (on qubits in the first case and on prime-dimensional qudits in the second case), which support transversal gates that are diagonal, non-Clifford and in the third level of the Clifford hierarchy. Calling this non-Clifford gate $U$ for convenience, they each make use of a distillation subroutine, which goes as follows, on either qubits or prime-dimensional qudits.
@@
    Prepare $k$ qubits/qudits in the state |$\ket{+}^{\otimes k}$, where $\ket{+}$ is an equal superposition of all computational basis states. Encode these qubits/qudits into the quantum code.
@@
Then, [13] and [19] concatenate this distillation subroutine to make a larger distillation routine that may arbitrarily suppress the error rate. In particular, the distillation routine will have multiple levels where the output of one level is used as the input to the next level. Where $[[n,k,d]]$ are the parameters of the quantum code at hand, both papers show that, in order to generate a supply of magic states with an error rate below $\epsilon$, the number of input magic states (of some (small) constant error rate) required per output state is
@@
which is called the overhead of the magic state distillation protocol, where the exponent $\gamma$ is [^13] Conventionally, magic state distillation is performed by concatenating protocols using an $[[n,k,d]]$ code $L$ times. The error rate becomes $\epsilon=(p_{\mathrm{ph}}/p_{\mathrm{th}})^{\mathcal{O}(d^{L})}$, i.e., $L\log d=\log(\mathcal{O}(\log(1/\epsilon)))$, where $p_{\mathrm{th}}$ is a threshold, and $p_{\mathrm{ph}}<p_{\mathrm{th}}$ is a physical error rate. The overhead is given by $(n/k)^{L}=\mathcal{O}(\log^{\gamma}(1/\epsilon))$ with $\gamma=\log(n/k)/\log d$. Note that this overhead scaling is not applicable to our protocol since we do not perform the concatenation but develop a way to suppress the error rate by a single-round protocol using asymptotically good triorthogonal codes.
@@ 6 Analysis of Error Suppression and Overhead
In this section, we analyse the error and the overhead of the distillation protocol presented in this paper. We consider the task of magic state distillation under the local stochastic error model described in Section 2.4. As described in Sections 4, our distillation protocol, using $[[n,k,d]]$ triorthogonal codes on $q$-dimensional qudits with $q=2^{s}$, starts with $\zeta=Cn$ noisy $CCZ$ states |$\ket{CCZ}$ at physical error rate $p_{\mathrm{ph}}$, converts them into $n$ noisy magic states |$\ket{M}$ on qudits, distils $k$ magic states |$\ket{M}$, and converts them into $l=k$ $CCZ$ states. In particular, as presented in Section 4, our protocol is designed for $s=10$, $C=O(1)$, $k=\Theta(n)$, and $d=\Theta(n)$. The efficient decoder of the algebraic geometry codes (i.e., of our triorthogonal codes) has the decoding radius of $t=\Theta(n)$ and thus can correct any $Z$-type errors (which are relevant for magic state distillation) up to weight $t$ on the $n$ qudits. Note that our protocol does not involve post-selection since we use the decoder for error correction rather than error detection (see also Section 4).
@@
Consider a family of protocols for magic state distillation using $[[n,k,d]]$ codes for $q$-dimensional qudits with decoding radius $t$, which starts with $\zeta=Cn$ noisy $CCZ$ states |$\ket{CCZ}$ at physical error rate $p_{\mathrm{ph}}$, converts them into $n$ noisy magic states |$\ket{M}$ on qudits, distils $k$ magic states |$\ket{M}$ without post-selection, and converts them into at least $\xi(\zeta)=k$ $CCZ$ states. Under the local stochastic error model, if the codes used in such a protocol satisfy
$k$ $=\Theta(n),$ (297)
$t$ $=\Theta(n),$ (298)
then the protocol has a constant threshold $p_{\mathrm{th}}>0$ such that if the physical error rate of the $\zeta$ initial $CCZ$ states is below the threshold, i.e., $p_{\mathrm{ph}}<p_{\mathrm{th}}$, then the protocol outputs $\xi(\zeta)$ final $CCZ$ states achieving a target error rate $\epsilon$ exponentially suppressed as
@@ Acknowledgements
All authors extend their thanks to N. Rengaswamy for discussions on CSS-T codes. A. Wills thanks The University of Tokyo, in particular the whole group of H. Yamasaki, for their funding of and hospitality during his visit to the institution, which led to this work. H. Yamasaki was supported by JST PRESTO Grant Number JPMJPR201A, JPMJPR23FC, JSPS KAKENHI Grant Number JP23K19970, and MEXT Quantum Leap Flagship Program (MEXT QLEAP) JPMXS0118069605, JPMXS0120351339.